# through $\mathcal{R}$eal-$\mathcal{T}$ime $\mathcal{V}$ideo

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) increasingly excel at perception, understanding, and reasoning. However, current benchmarks inadequately evaluate their ability to perform these tasks continuously in dynamic, real-world environments. To bridge this gap, we introduce $\mathcal{RTV}\text{-}Bench$, a fine-grained benchmark for MLLM real-time video analysis. $\mathcal{RTV}\text{-}Bench$ includes three key principles: (1) Multi-Timestamp Question Answering (MTQA), where answers evolve with scene changes; (2) Hierarchical Question Structure, combining basic and advanced queries; and (3) Multi-dimensional Evaluation, assessing the ability of continuous perception, understanding, and reasoning. $\mathcal{RTV}\text{-}Bench$ contains 552 diverse videos (167.2 hours) and 4,631 high-quality QA pairs. We evaluated leading MLLMs, including proprietary (GPT-4o, Gemini 2.0), open-source offline (Qwen2.5-VL, VideoLLaMA3), and open-source real-time (VITA-1.5, InternLM-XComposer2.5-OmniLive) models. Experiment results show open-source real-time models largely outperform offline ones but still trail top proprietary models. Our analysis also reveals that larger model size or higher frame sampling rates do not significantly boost $\mathcal{RTV}\text{-}Bench$ performance, sometimes causing slight decreases. This underscores the need for better model architectures optimized for video stream processing and long sequences to advance real-time video analysis with MLLMs. Our benchmark toolkit is available at Project Page.
**footnotetext: Both authors made equal core contributions.$\dagger$\$$\dagger$\$footnotetext: Corresponding author.
@@ 1 Introduction
The ability to comprehend and respond to complex real-world scenarios in real-time represents a crucial challenge in the pursuit of general artificial intelligence Dang et al. (2024); Zhang et al. (2025b); Hu et al. (2025). Drawing inspiration from the remarkable achievements of large language models (LLMs) in a wide range of tasks Liu et al. (2024); Huang et al. (2025), the potential of multimodal large language models (MLLMs) is increasingly being realized in the domain of visual scene analysis. The research trajectory of Video-LLMs Lin et al. (2023); Maaz et al. (2024); Cheng et al. (2024); Wang et al. (2024a) is progressively advancing from the understanding of short, vision-centric videos Li et al. (2023a); Jin et al. (2024); Lin et al. (2024a); Li et al. (2024d) to the integration of omni-modal information—including video, audio, and subtitles Chen et al. (2023); Duan et al. (2023); Li et al. (2023b); Tan et al. (2024); Li et al. (2024c)—for comprehensive long video analysis.
@@
Recently, VStream Zhang et al. (2024a) was the first to attempt to test this capability, with a focus primarily on extending the duration of videos. In addition, both StreamingBench Xiong et al. (2025) and OVOBench Li et al. (2025) have made varying degrees of improvements in the types and standards of assessment. However, their evaluation of real-time responsiveness is often inadequate, overlooking the capacity to capture transitions and fleeting details from visual input that arrive sequentially, not instantaneously. This limitation highlights the need for a more focused assessment of continuous analysis abilities.
@@
Based on the above considerations, we introduce $\mathcal{RTV}\text{-}Bench$, featuring three core innovations designed to benchmark the continuous analysis capabilities—specifically perception, understanding, and reasoning—of MLLMs within real-time video contexts. First, the Multi-Timestamp Q&A mechanism challenges real-time tracking and state update by posing queries whose answers evolve within a video. Crucially, unlike benchmarks like OVO-Bench Li et al. (2025) that typically introduce different questions at different timestamps, RTV-Bench revisits the same conceptual query, where only the correct answer shifts as the scene unfolds. This approach more rigorously tests the model capacities for continuous analysis in real-time scenarios, surpassing single-query-single-answer and static evaluations. Second, the Hierarchical Question Structure enforces reliable, sequential reasoning by employing a basic-to-advanced design where higher-order questions logically depend on grasping foundational perceptions and understanding, thus mitigating cognitive shortcuts. Finally, Multidimensional Evaluation moves beyond single scores to provide granular diagnostics, assessing performance across eight fine-grained dimensions critical to continuous analysis in dynamic scenarios, offering a more insightful perspective on model capabilities and bottlenecks in handling real-time video information, detailed cases are shown in Figure 1.
@@
Through RTV-Bench’s novel design and the detailed evaluations presented herein, we provide comprehensive insights into the current state of MLLMs for continuous video analysis: Results for all tested models on RTV-Bench reveal significant bottlenecks for existing MLLMs in real-time video analysis, as (1) most models achieve accuracy below 50%. Furthermore, (2) model performance shows minimal correlation with the number of input frames or model size. Notably, (3) models specifically designed for real-time video significantly outperform traditional offline video models; even the lowest-performing real-time model tested (VITA-1.5 Fu et al. (2025)) surpasses a traditional offline counterpart (VideoLLaMA2 Cheng et al. (2024)). Building upon these findings, we further discuss promising research trajectories to enhance MLLM online video processing. We expect that our benchmark, analysis, and outlined future directions will contribute significantly to the development of more robust and capable online video MLLMs.
@@ 2 RTV-Bench
2.1 Challenge for Video Analysis
Existing benchmarks for long video analysis Fu et al. (2024); Wu et al. (2025) attempt to gauge the true capacity of MLLMs by extending video duration, enhancing the difficulty of QA tasks, and introducing new sub-tasks. Other works Lin et al. (2024b) focus on real-time video scenarios, evaluating model responsiveness to queries in real-time contexts. However, current models face the risk of memory loss and attention shift, both in long and real-time video contexts. We categorize these challenges as failures in the continuous analysis capability of MLLMs. In video scenarios, MLLMs are considered to possess strong continuous analysis ability only if they can effectively recall prior visual information to accurately respond to user queries (similar to needle-in-a-haystack tasks in natural language settings) and maintain robust perception of incoming streams of new visual data and queries (akin to multimodal multi-turn dialogues).
@@ 2.2 Benchmark Overview
The RTV-Bench is designed to assess a model’s ability to perceive, understand, and reason in long video and real-time streaming contexts. Models should be adept at recognizing correct temporal query patterns. For example, in an ongoing sports match, when a user inquires about the goalkeeper’s actions, the model should respond with the situation at the query time, rather than historical moments. However, it should also retrieve relevant information from memory, such as identifying the goalkeeper from earlier footage. This necessitates that models not only possess exceptionally high fundamental scene comprehension capabilities, but also demonstrate effective understanding of all stages within the temporal flow—namely, comprehensive spatiotemporal comprehension encompassing the past, present, and future. To this end, we propose a series of sub-tasks for video continuous analysis, addressing both spatiotemporal elements within the video and the intrinsic capabilities of the model. The intrinsic model capabilities are divided into eight categories: Temporal Perception (TP), Scene Perception (SP), Visual Perception (VP), Future Prediction (FP), Phenomenological Understanding (PU), Intent Analysis (IA), Global Understanding (GU), and Spatiotemporal Reasoning (SR), see Figure 1 for details.
@@ 2.3 Benchmark Construction
Benchmark MT Num. of QA Avg. Dur. (min) Total Dur. (h) Labels
Video-MME Fu et al. (2024) 2.7k 17 254.0 M
MSRVTT Xu et al. (2016) 73k 0.25 12.5 A
MSVD-QA Chen and Dolan (2011) 13k 0.17 1.4 A
MovieChat-1k Song et al. (2024) 13k 9.4 156.0 M
MVBench Li et al. (2024b) 4k 0.25 34.5 A+M
ActivityNet-QA Yu et al. (2019) 58k 1.87 25.0 M
Vstream-Q Zhang et al. (2024a) 3.5k 40 21.0
StreamBench Wu et al. (2024) 1.8k 4.5 25.0 M
OVO-Bench Li et al. (2025) 2.8k 6 66.0 A+M
RTV-Bench (Ours) 4.6k 18 167.2 M
@@
Our video data sources include EgoSchema Mangalam et al. (2023) and publicly available online videos. Unlike most existing benchmarks, we incorporated manual review during the collection phase. Three data collectors sourced videos from various domains and manually excluded highly similar videos, focusing on long videos with high dynamics and real-time needs. These sources ensure targeted and scientific evaluation of video models’ real-time capabilities.
@@
Rigorous manual annotation by qualified experts was essential for RTV-Bench’s reliability in assessing continuous video analysis. While an LLM (Deepseek Liu et al. (2024)) provided initial question templates, human annotators significantly refined questions to target dynamic scenes and temporal reasoning. They meticulously crafted questions where answers change over time and identified the earliest valid timestamp for each answer option in MTQA. This human-centric, multi-annotator process ensured robustness and created evaluations specifically focused on temporal dynamics in video.
@@ 3.1 Experiment Setup
Our experiments were conducted across two A800 GPUs to comprehensively evaluate the performance of mainstream MLLMs on our benchmark. The selected models include VideoLLaMA2 Cheng et al. (2024), VideoLLaMA3 Zhang et al. (2025a), GPT-4O Hurst et al. (2024), InternLM-XComposer2.5-OmniLive (IXC2.5-OL) Zhang et al. (2024b), VITA1.5 Fu et al. (2025), LLaVA-Video Lin et al. (2023), LLaVA-OneVision Li et al. (2024a), and Qwen2.5-VL Yang et al. (2024). These models were primarily tested with 7 billion parameters (7B) to ensure a fair comparison under similar computational constraints.
@@
We compare two model categories: traditional offline models ($M_{offline}$) and novel real-time online models ($M_{online}$). These categories differ significantly in architecture, training, and data requirements. Architecturally, $M_{offline}$ typically employ sequential vision encoder-decoders, often limited by fixed context windows and incurring higher processing latency. In contrast, $M_{online}$ are designed for continuous, low-latency ingestion of the video stream $V$. They prioritize maintaining an internal state $S_{t}$ that summarizes the video information processed up to time $t$ (i.e., $V[0,t]$), enabling real-time responsiveness. Models like IXC2.5-OL Zhang et al. (2024b) implement this using techniques like modular parallelism and dedicated long-term memory. These architectural distinctions lead to different training paradigms and data needs. $M_{offline}$ usually rely on end-to-end fine-tuning using standard annotated video datasets. $M_{online}$, however, often require specialized training strategies (e.g., VITA-1.5’s staged fusion, IXC2.5-OL’s targeted training for memory and interaction) and benefit most from specialized corpora designed for long-duration, interactive streaming scenarios. The ability of $M_{online}$ to maintain and update state $S_{t}$ is particularly relevant for RTV-Bench’s Multi-Timestamp QA (MTQA) challenge, where the correct answer $A^{*}(Q,t_{q})$ to a query $Q$ depends on the specific query time $t_{q}$. To evaluate both model types on this benchmark, we adapt the testing procedure. For $M_{online}$, queries $Q$ are presented at their timestamp $t_{q}$, and the models leverage their continuously updated state $S_{t_{q}}$ to generate the answer $A_{M_{online}}=M_{online}(Q,t_{q}|S_{t_{q}})$. Since $M_{offline}$ lack this inherent streaming capability, we simulate real-time interaction for them: when a query $Q_{i}$ is posed at time $t_{q,i}$, we extract and provide only the relevant video segment $V_{i}$ (corresponding to the query’s context). The offline model’s answer is thus based solely on this isolated segment: $A_{M_{offline},i}=M_{offline}(Q_{i},V_{i})$.
@@
Qwen2.5-VL 7B 33.53/3.53 31.39/3.70 33.65/5.20 31.37 34.37 32.75/4.10
VideoLLaMA2 7B 40.62/8.67 39.85/7.77 37.49/6.75 45.77 34.95 39.55/7.90
VideoLLaMA3 7B 37.98/5.83 35.29/5.73 35.78/6.80 38.62 34.91 36.42/6.10
LLaVA-OneVision 7B 35.38/3.97 34.21/4.63 33.57/4.95 35.80 33.58 34.49/4.40
LLaVA-Video 7B 35.83/5.03 33.81/3.77 35.15/5.75 36.28 34.17 34.90/4.80
@@
GPT-4o 51.61/21.90 49.31/20.76 48.71/23.95 56.53 44.73 50.02/22.10
Gemini 2.0 Flash 41.67/11.00 42.71/12.73 41.44/12.05 47.49 38.64 42.00/12.00
@@
Qwen2.5-VL 7B 32.37 / 2.4 37.48 / 5.8 30.73 / 2.4 33.53 / 3.53
VideoLLaMA2 7B 39.52 / 7.9 42.49 / 9.4 39.85 / 8.7 40.62 / 8.67
VideoLLaMA3 7B 37.82 / 6.1 39.24 / 7.6 36.87 / 3.8 37.98 / 5.83
LLaVA-OneVision 7B 35.09 / 3.9 35.86 / 3.8 35.20 / 4.2 35.38 / 3.97
LLaVA-Video 7B 34.07 / 4.8 38.97 / 5.8 34.45 / 4.5 35.83 / 5.03
@@
GPT-4o 48.60 / 18.2 53.59 / 23.4 52.63 / 24.1 51.61 / 21.90
Gemini 2.0 Flash 40.49 / 9.5 45.19 / 16.1 39.34 / 7.4 41.67 / 11.00
Table 3: Detailed evaluation results on the category of Perception. Temporal Perception (TP), Visual Perception (VP) and Scene Perception (SP)
@@
Qwen2.5-VL 7B 29.11 / 2.2 35.69 / 6.9 29.36 / 2.0 31.39 / 3.70
VideoLLaMA2 7B 37.34 / 7.6 42.21 / 9.5 40.92 / 6.2 39.85 / 7.77
VideoLLaMA3 7B 33.54 / 5.8 39.13 / 5.9 33.39 / 4.3 35.35 / 5.33
LLaVA-OneVision 7B 32.07 / 4.3 33.51 / 3.0 37.06 / 6.6 34.21 / 4.63
LLaVA-Video 7B 29.42 / 2.5 35.69 / 3.9 36.33 / 4.9 33.81 / 3.77
@@
GPT-4o 45.02 / 15.7 54.32 / 25.8 48.58 / 20.8 49.31 / 20.76
Gemini 2.0 Flash 35.70 / 10.6 45.65 / 11.3 46.78 / 16.3 42.71 / 12.73
Table 4: Detailed evaluation results on the category of Understanding. Phenomenological Understanding (PU), Global Understanding (GU) and Intent Analysis (IA)
@@
Qwen2.5-VL 7B 35.33 / 4.4 33.67 / 4.5 34.50 / 4.45
VideoLLaMA2 7B 41.47 / 7.5 33.50 / 6.0 37.49 / 6.75
VideoLLaMA3 7B 38.05 / 6.9 33.84 / 3.9 35.95 / 5.40
LLaVA-OneVision 7B 38.23 / 7.2 28.91 / 2.7 33.57 / 4.95
LLaVA-Video 7B 39.08 / 9.1 31.22 / 2.4 35.15 / 5.75
@@
GPT-4o 54.67 / 27.1 42.75 / 20.8 48.71 / 23.95
Gemini 2.0 Flash 44.42 / 13.6 38.46 / 10.5 41.44 / 12.05
3.2 Experiment Results
Online vs. Offline Models.
RTV-Bench results (shown in Table 2) demonstrate that online models designed for real-time processing, particularly IXC2.5-OL, outperform offline models in overall performance (e.g., IXC2.5-OL: 47.33% Acc / 15.40 Score vs. VideoLLaMA2: 39.55% Acc / 7.90 Score). This advantage is notable in tasks requiring temporal analysis, such as Multi-Timestamp Question Answering (MTQA), where online models like IXC2.5-OL (38.21% Acc) surpass their offline counterparts (typically 33-35% Acc). We observe that (Table 4) all models excel at Visual Perception tasks, likely because these tasks are more image-centric compared to Temporal or Scene Perception, and current video models often build upon image architectures. This suggests a need for more specialized video model designs, such as dedicated spatio-temporal architectures or real-time processing capabilities.
@@ 3.2 Experiment Results
However, a significant gap remains compared to leading closed-source models like GPT-4o. GPT-4o exhibits superior performance across the board, especially in Overall Score (22.10) and on complex tasks like Understanding and Reasoning (Tables 4 and 5). Evidently, top-tier closed-source models still hold a significant edge, particularly in complex cognitive capabilities.
Accuracy vs. Score.
The Score metric is crucial for revealing reliable, hierarchical reasoning beyond simple Accuracy. While GPT-4o’s Accuracy gain over IXC2.5-OL is moderate (50.02% vs. 47.33%), its substantially higher Score (22.10 vs. 15.40) indicates a more robust reasoning process, reliably building upon foundational understanding to address complex queries. Conversely, lower Scores relative to Accuracy, common among open-source models, suggest instability in multi-step reasoning. Thus, as shown in Figure 3.the Score metric effectively quantifies the deeper, more reliable comprehension capabilities demonstrated by models like GPT-4o on RTV-Bench’s challenging tasks and it can distinguish model performance more clearly.
@@
To investigate the effect of model scale, we evaluated Qwen2.5-VL with different parameter sizes. As depicted in Figure 3, the performance across different model sizes exhibits limited variation. Accuracy remains relatively consistent across the tested configurations. For instance, Overall Accuracy hovers narrowly between approximately 31.4% and 32.8%. The limited benefit from scaling suggests the core challenge for real-time performance lies less in model size and more in model designs. Larger models failing to improve significantly implies a potential bottleneck in how effectively their internal representations capture the demands of real-time processing.

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) increasingly excel at perception, understanding, and reasoning. However, current benchmarks inadequately evaluate their ability to perform these tasks continuously in dynamic, real-world environments. To bridge this gap, we introduce $\mathcal{RTV}\text{-}Bench$, a fine-grained benchmark for MLLM real-time video analysis. $\mathcal{RTV}\text{-}Bench$ includes three key principles: (1) Multi-Timestamp Question Answering (MTQA), where answers evolve with scene changes; (2) Hierarchical Question Structure, combining basic and advanced queries; and (3) Multi-dimensional Evaluation, assessing the ability of continuous perception, understanding, and reasoning. $\mathcal{RTV}\text{-}Bench$ contains 552 diverse videos (167.2 hours) and 4,631 high-quality QA pairs. We evaluated leading MLLMs, including proprietary (GPT-4o, Gemini 2.0), open-source offline (Qwen2.5-VL, VideoLLaMA3), and open-source real-time (VITA-1.5, InternLM-XComposer2.5-OmniLive) models. Experiment results show open-source real-time models largely outperform offline ones but still trail top proprietary models. Our analysis also reveals that larger model size or higher frame sampling rates do not significantly boost $\mathcal{RTV}\text{-}Bench$ performance, sometimes causing slight decreases. This underscores the need for better model architectures optimized for video stream processing and long sequences to advance real-time video analysis with MLLMs.
@@ 1 Introduction
The ability to comprehend and respond to complex real-world scenarios in real-time represents a crucial challenge in the pursuit of general artificial intelligence [6, 24, 10]. Drawing inspiration from the remarkable achievements of large language models (LLMs) in a wide range of tasks [25, 11, 2], the potential of multimodal large language models (MLLMs) is increasingly being realized in the domain of visual scene analysis [48, 41, 5]. The research trajectory of Video-LLMs [21, 27, 4, 35, 33] is progressively advancing from the understanding of short, vision-centric videos [15, 13, 22, 19] to the integration of omni-modal information—including video, audio, and subtitles [3, 7, 17, 32, 18, 26]—for comprehensive long video analysis.
@@
Recently, VStream [45] was the first to attempt to test this capability, with a focus primarily on extending the duration of videos. In addition, both StreamingBench [39] and OVOBench [20] have made varying degrees of improvements in the types and standards of assessment. However, their evaluation of real-time responsiveness is often inadequate, overlooking the capacity to capture transitions and fleeting details from visual input that arrive sequentially, not instantaneously. This limitation highlights the need for a more focused assessment of continuous analysis abilities.
@@
Based on the above considerations, we introduce $\mathcal{RTV}\text{-}Bench$, featuring three core innovations designed to benchmark the continuous analysis capabilities—specifically perception, understanding, and reasoning—of MLLMs within real-time video contexts. First, the Multi-Timestamp Q&A mechanism challenges real-time tracking and state update by posing queries whose answers evolve within a video. Crucially, unlike benchmarks like OVO-Bench [20] that typically introduce different questions at different timestamps, RTV-Bench revisits the same conceptual query, where only the correct answer shifts as the scene unfolds. This approach more rigorously tests the model capacities for continuous analysis in real-time scenarios, surpassing single-query-single-answer and static evaluations. Second, the Hierarchical Question Structure enforces reliable, sequential reasoning by employing a basic-to-advanced design where higher-order questions logically depend on grasping foundational perceptions and understanding, thus mitigating cognitive shortcuts. Finally, Multidimensional Evaluation moves beyond single scores to provide granular diagnostics, assessing performance across eight fine-grained dimensions critical to continuous analysis in dynamic scenarios, offering a more insightful perspective on model capabilities and bottlenecks in handling real-time video information, detailed cases are shown in Figure 1.
@@
Through RTV-Bench’s novel design and the detailed evaluations presented herein, we provide comprehensive insights into the current state of MLLMs for continuous video analysis: Results for all tested models on RTV-Bench reveal significant bottlenecks for existing MLLMs in real-time video analysis, as (1) most models achieve accuracy below 50%. Furthermore, (2) model performance shows minimal correlation with the number of input frames or model size. Notably, (3) models specifically designed for real-time video significantly outperform traditional offline video models, even the lowest-performing real-time model tested (VITA-1.5 [9]) surpasses a traditional offline counterpart (VideoLLaMA2 [4]). Building upon these findings, we further discuss promising research trajectories to enhance MLLM online video processing. We expect that our benchmark, analysis, and outlined future directions will contribute significantly to the development of more robust and capable online video MLLMs.
@@ 2 RTV-Bench
2.1 Challenge for Real-Time Video Analysis
Existing benchmarks for long video analysis [8, 38] attempt to gauge the true capacity of MLLMs by extending video duration, enhancing the difficulty of QA tasks, and introducing new sub-tasks. Other works [23] focus on real-time video scenarios, evaluating model responsiveness to queries in real-time contexts. However, current models face the risk of memory loss and attention shift, both in long and real-time video contexts. We categorize these challenges as failures in the continuous analysis capability of MLLMs. In video scenarios, MLLMs are considered to possess strong continuous analysis ability only if they can effectively recall prior visual information to accurately respond to user queries (similar to needle-in-a-haystack tasks in natural language settings) and maintain robust perception of incoming streams of new visual data and queries (akin to multimodal multi-turn dialogues).
@@ 2.2 Benchmark Overview
The RTV-Bench is designed to assess a model’s ability to perceive, understand, and reason in long video and real-time streaming contexts. Models should be adept at recognizing correct temporal query patterns. For example, during live sports queries (e.g., goalkeeper actions), models must prioritize current contexts over historical data. However, it should also retrieve relevant information from memory, such as identifying the goalkeeper from earlier footage. This necessitates that models not only possess exceptionally high fundamental scene comprehension capabilities, but also demonstrate effective understanding of all stages within the temporal flow—namely, comprehensive spatiotemporal comprehension encompassing the past, present, and future. To this end, we propose a series of sub-tasks for video continuous analysis, addressing both spatiotemporal elements within the video and the intrinsic capabilities of the model. The intrinsic model capabilities are divided into eight categories: Temporal Perception (TP), Scene Perception (SP), Visual Perception (VP), Future Prediction (FP), Phenomenological Understanding (PU), Intent Analysis (IA), Global Understanding (GU), and Spatiotemporal Reasoning (SR), see Figure 1 for details.
@@ 2.3 Benchmark Construction
Our video data sources include EgoSchema [28] and publicly available online videos. Unlike most existing benchmarks, we incorporated manual review during the collection phase. Three data collectors sourced videos from various domains and manually excluded highly similar videos, focusing on long videos with high dynamics and real-time needs. These sources ensure targeted and scientific evaluation of video models’ real-time capabilities.
Benchmark MT Num. of QA Avg. Dur. (min) Total Dur. (h) Labels
Video-MME [8] 2.7k 17 254.0 M
MSRVTT [40] 73k 0.25 12.5 A
MSVD-QA [1] 13k 0.17 1.4 A
MovieChat-1k [31] 13k 9.4 156.0 M
MVBench [16] 4k 0.25 34.5 A+M
ActivityNet-QA [43] 58k 1.87 25.0 M
Vstream-Q [45] 3.5k 40 21.0
StreamBench [37] 1.8k 4.5 25.0 M
OVO-Bench [20] 2.8k 6 66.0 A+M
RTV-Bench (Ours) 4.6k 18 167.2 M
@@
Rigorous manual annotation by qualified experts was essential for RTV-Bench’s reliability in assessing continuous video analysis. While an LLM (Deepseek [25]) provided initial question templates, human annotators significantly refined questions to target dynamic scenes and temporal reasoning. They meticulously crafted questions where answers change over time and identified the earliest valid timestamp for each answer option in MTQA. This human-centric, multi-annotator process ensured robustness and created evaluations specifically focused on temporal dynamics in video.
@@ 3.1 Experiment Setup
Our experiments were conducted across two A800 GPUs to comprehensively evaluate the performance of mainstream MLLMs on our benchmark. The selected models include VideoLLaMA2 [4], VideoLLaMA3 [44], GPT-4O [12], InternLM-XComposer2.5-OmniLive (IXC2.5-OL) [46], VITA1.5 [9], LLaVA-Video [21], LLaVA-OneVision [14], and Qwen2.5-VL [42]. These models were primarily tested with 7B to ensure a fair comparison under similar computational constraints.
@@
We compare two model categories: traditional offline models ($M_{\text{offline}}$) and novel real-time online models ($M_{\text{online}}$). These categories differ significantly in architecture, training, and data requirements. Architecturally, $M_{\text{offline}}$ typically employ sequential vision encoder-decoders, often limited by fixed context windows and incurring higher processing latency. In contrast, $M_{\text{online}}$ are designed for continuous, low-latency ingestion of the video stream $V$. They prioritize maintaining an internal state $S_{t}$ that summarizes the video information processed up to time $t$ (i.e., $V[0,t]$), enabling real-time responsiveness. Models like IXC2.5-OL [46] implement this using techniques like modular parallelism and dedicated long-term memory. These architectural distinctions lead to different training paradigms and data needs. $M_{\text{offline}}$ usually rely on end-to-end fine-tuning using standard annotated video datasets. $M_{\text{online}}$, however, often require specialized training strategies (e.g., VITA-1.5’s staged fusion, IXC2.5-OL’s targeted training for memory and interaction) and benefit most from specialized corpora designed for long-duration, interactive streaming scenarios. The ability of $M_{\text{online}}$ to maintain and update state $S_{t}$ is particularly relevant for RTV-Bench’s Multi-Timestamp QA (MTQA) challenge, where the correct answer $A^{*}(Q,t_{q})$ to a query $Q$ depends on the specific query time $t_{q}$. To evaluate both model types on this benchmark, we adapt the testing procedure. For $M_{\text{online}}$, queries $Q$ are presented at their timestamp $t_{q}$, and the models leverage their continuously updated state $S_{t_{q}}$ to generate the answer $A_{M_{\text{online}}}=M_{\text{online}}(Q,t_{q}|S_{t_{q}})$. Since $M_{\text{offline}}$ lack this inherent streaming capability, we simulate real-time interaction for them: when a query $Q_{i}$ is posed at time $t_{q,i}$, we extract and provide only the relevant video segment $V_{i}$ (corresponding to the query’s context). The offline model’s answer is thus based solely on this isolated segment: $A_{M_{\text{offline}},i}=M_{\text{offline}}(Q_{i},V_{i})$.
@@
Qwen2.5-VL [42] 7B 33.53/3.53 31.39/3.70 33.65/5.20 31.37 34.37 32.75/4.10
VideoLLaMA2 [4] 7B 40.62/8.67 39.85/7.77 37.49/6.75 45.77 34.95 39.55/7.90
VideoLLaMA3 [44] 7B 37.98/5.83 35.29/5.73 35.78/6.80 38.62 34.91 36.42/6.10
LLaVA-OneVision [14] 7B 35.38/3.97 34.21/4.63 33.57/4.95 35.80 33.58 34.49/4.40
LLaVA-Video [21] 7B 35.83/5.03 33.81/3.77 35.15/5.75 36.28 34.17 34.90/4.80
@@
GPT-4o [12] 51.61/21.90 49.31/20.76 48.71/23.95 56.53 44.73 50.02/22.10
Gemini 2.0 Flash [34] 41.67/11.00 42.71/12.73 41.44/12.05 47.49 38.64 42.00/12.00
3.2 Experiment Results
Online vs. Offline Models.
As shown in Table 2, online models optimized for real-time processing—particularly IXC2.5-OL—surpass offline counterparts in overall performance metrics. IXC2.5-OL achieves 47.33% Accuracy and 15.40 Score, significantly outperforming offline models like VideoLLaMA2 (39.55% Accuracy / 7.90 Score). Furthermore, when comparing online models, IXC2.5-OL demonstrates a clear advantage over VITA-1.5 with improvements of 2.82% in Accuracy and 3.6 points in Score. A notable performance gap emerges in temporal analysis tasks: IXC2.5-OL attains 38.21% Accuracy in Multi-Timestamp Question Answering (MTQA), substantially higher than the 33–35% range typical of offline models. This notable discrepancy suggests that current online models may be promising avenues toward continuous analysis capabilities.
Table 3: Detailed evaluation results on the category of Perception. Temporal Perception (TP), Visual Perception (VP) and Scene Perception (SP).
@@ 3.2 Experiment Results
Qwen2.5-VL [42] 7B 32.37 / 2.4 37.48 / 5.8 30.73 / 2.4 33.53 / 3.53
VideoLLaMA2 [4] 7B 39.52 / 7.9 42.49 / 9.4 39.85 / 8.7 40.62 / 8.67
VideoLLaMA3 [44] 7B 37.82 / 6.1 39.24 / 7.6 36.87 / 3.8 37.98 / 5.83
LLaVA-OneVision [14] 7B 35.09 / 3.9 35.86 / 3.8 35.20 / 4.2 35.38 / 3.97
LLaVA-Video [21] 7B 34.07 / 4.8 38.97 / 5.8 34.45 / 4.5 35.83 / 5.03
@@
GPT-4o [12] 48.60 / 18.2 53.59 / 23.4 52.63 / 24.1 51.61 / 21.90
Gemini 2.0 Flash [34] 40.49 / 9.5 45.19 / 16.1 39.34 / 7.4 41.67 / 11.00
Table 4: Detailed evaluation results on the category of Understanding. Phenomenological Understanding (PU), Global Understanding (GU) and Intent Analysis (IA).
@@
Qwen2.5-VL [42] 7B 29.11 / 2.2 35.69 / 6.9 29.36 / 2.0 31.39 / 3.70
VideoLLaMA2 [4] 7B 37.34 / 7.6 42.21 / 9.5 40.92 / 6.2 39.85 / 7.77
VideoLLaMA3 [44] 7B 33.54 / 5.8 39.13 / 5.9 33.39 / 4.3 35.35 / 5.33
LLaVA-OneVision [14] 7B 32.07 / 4.3 33.51 / 3.0 37.06 / 6.6 34.21 / 4.63
LLaVA-Video [21] 7B 29.42 / 2.5 35.69 / 3.9 36.33 / 4.9 33.81 / 3.77
@@
GPT-4o [12] 45.02 / 15.7 54.32 / 25.8 48.58 / 20.8 49.31 / 20.76
Gemini 2.0 Flash [34] 35.70 / 10.6 45.65 / 11.3 46.78 / 16.3 42.71 / 12.73
@@
Qwen2.5-VL [42] 7B 35.33 / 4.4 33.67 / 4.5 34.50 / 4.45
VideoLLaMA2 [4] 7B 41.47 / 7.5 33.50 / 6.0 37.49 / 6.75
VideoLLaMA3 [44] 7B 38.05 / 6.9 33.84 / 3.9 35.95 / 5.40
LLaVA-OneVision [14] 7B 38.23 / 7.2 28.91 / 2.7 33.57 / 4.95
LLaVA-Video [21] 7B 39.08 / 9.1 31.22 / 2.4 35.15 / 5.75
@@
GPT-4o [12] 54.67 / 27.1 42.75 / 20.8 48.71 / 23.95
Gemini [34] 2.0 Flash 44.42 / 13.6 38.46 / 10.5 41.44 / 12.05
@@
While a performance gap persists compared to leading closed-source models like GPT-4o, state-of-the-art online architectures demonstrate remarkable progress. The online model IXC2.5-OL achieves near-top-tier performance with 47.33% Overall Accuracy and 15.40 Score, substantially outperforming mid-range closed-source systems like Gemini 2.0 Flash. Notably, IXC2.5-OL closes the accuracy gap with GPT-4o to 4.4% in perception tasks(Tables 4) and 1.1% in video understanding, demonstrating competitive performance in multimodal domains. However, limitations emerge in complex reasoning where GPT-4o maintains decisive advantages, especially on complex tasks like Understanding and Reasoning(Tables 4 and 5). This pattern highlights that while modern online models have approached entry-level commercial systems and even challenged premium models in specific competencies, structural innovations remain critical to bridge gaps in advanced cognitive tasks like multi-step reasoning and temporal analysis.
@@
To investigate the effect of model scale, we evaluated Qwen2.5-VL across 3B-72B parameters. Figure 3 reveals non-monotonic scaling: medium-sized models (7B-34B) outperform both smaller and larger variants. All metrics fluctuate within a 5% range, with no configuration exceeding 34B’s maximums. This scale saturation implies current model architectures may lack the representational efficiency required for effective real-time processing, emphasizing the need for structural innovations rather than pure parameter growth.